Prove that (Tanx)/secx+1 = (secx-1)/tanx using trigonometric identities and where x is an angle.
Answer by Alan3354(69443)

Prove that (Tanx)/secx+1 = (secx-1)/tanx using trigonometric identities and where x is an angle.
------------
(Tanx)/secx+1 = (secx-1)/tanx
Cross multiply
tan^2 = sec^2 - 1
tan^2 + 1 = sec^2
I'm convinced.
-----
If someone can show an example where working one side only makes a difference I'd like to see it.